Jeffrey Mitton and his colleagues found three genotypes(R2R2,R2R3 and R3R3)at a locus encoding the enzyme peroxidase in Penderosa pine tree growing at Glacier Lake,Colorado.The observed number of genotypes were:

Genotype

R2R2

R2R3

R3R3

Number observed

135

44

11

Using CHI square test,determine whether the ponderosa pine tree at Glacier Lake conform to be Hardy-Weinberg equilibrium at the peroxidase locus?


1
Expert's answer
2021-09-13T17:11:02-0400

R2R2 P2 = 135

R2R3 2Pq = 44

R3R3 q2 = 11

Total = 190

R3R3 or q2 = 11 /190 = 0.05

Allelic frequency R3 or q = 0.22

P + q = 1

P = 1-0.22 = 0.78

P2 = 0.78 X 0.78 = 0.60 2pq = 2 X .78 X .22 = 0.34 q2 = .22 X .22 = 0.04


Chi square test


Degree of freedom 1 at 95% chi square table value = 3.814

12.12>3.814

we conclude that the genotype frequencies in this population are significantly different than what would be expected so this is not Hardy-Weinberg equilibrium.